# Do Existing Testing Tools Really Uncover Gender Bias in Text-to-Image Models?

### v1

@@ Abstract.
Text-to-Image (T2I) models have recently gained significant attention due to their ability to generate high-quality images and are consequently used in a wide range of applications. However, there are concerns about the gender bias of these models. Previous studies have shown that T2I models can perpetuate or even amplify gender stereotypes when provided with neutral text prompts (e.g., ‘a photo of a CEO’ is often associates with male images, while ‘a photo of nurse’ is often associates with female images). Researchers have proposed automated gender bias uncovering detectors for T2I models, but a crucial gap exists: no existing work comprehensively compares the various detectors and understands how the gender bias detected by them deviates from the actual situation. This study addresses this gap by validating previous gender bias detectors using a manually labeled dataset and comparing how the bias identified by various detectors deviates from the actual bias in T2I models, as verified by manual confirmation. We create a dataset consisting of 6,000 images generated from three cutting-edge T2I models, Stable Diffusion XL, Stable Diffusion 3, and Dreamlike Photoreal 2.0. During the human-labeling process, we find that all three T2I models generate a portion (12.48% on average) of low-quality images (e.g., generate images with no face present), where human annotators cannot determine the gender of the person. Our analysis reveals that all three T2I models show a preference for generating male images, with SDXL being the most biased. Additionally, images generated using prompts containing professional descriptions (e.g., lawyer or doctor) show the most bias. We evaluate seven gender bias detectors and find that none fully capture the actual level of bias in T2I models, with some detectors overestimating bias by up to 26.95%. We further investigate the causes of inaccurate estimations, highlighting the limitations of detectors in dealing with low-quality images. Based on our findings, we propose an enhanced detector called CLIP-Enhance, which most accurately measures the gender bias in T2I models, with a difference of only 0.47%-1.23%, and most effectively filters out 82.91% of low-quality images. [^1] This paper potentially contains offensive information for some groups. We have made our dataset and code publicly available. [^2] https://doi.org/10.6084/m9.figshare.27377649.v1
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Despite the advanced capabilities of T2I models, concerns about gender bias that are potentially demonstrated by these models remain significant. Previous studies have shown that T2I models tend to associate males with descriptions of high-paying jobs, such as CEO, lawyer, and doctor, while associating females with low-paying jobs, such as housekeeper and cashier (Cho et al., 2023; Nicoletti and Bass, 2023). Bloomberg further highlighted how T2I models can exacerbate gender bias (Nicoletti and Bass, 2023). For example, women made up only about 3% of the images generated by the T2I model (Stable Diffusion v1.5) for the keyword “judge,” despite the fact that 34% of US judges are women. As AI-generated images increasingly permeate our daily lives (Name, 2024), the stereotypes they reinforce and the neglect of minority genders may deepen existing unfairness. Therefore, it is crucial to uncover gender bias in T2I models.
@@
  • Detector Enhancement. We enhance the detector, achieving the best performance in accurately reflecting the gender bias in T2I models with a difference from the ground truth of only 0.47%-1.23% and effectively filtering out 82.91% of low-quality images.
@@
The rest of this paper is organized as follows. Section 2 discusses the background of our study. Section 3 presents the design of our study. Section 4 shows the evaluation results and answers to our research questions. Section 6 discusses the related work. Threats to validity are discussed in Section 7. Finally, we conclude our work and present future directions in Section 8.
@@ 2.1. Text-to-Image Models
A T2I model takes a textual description as input and generates a corresponding image that matches the description as output. With the advancement of deep learning techniques, T2I generation models began to evolve in the mid-2010s. Initially, Variational Auto Encoders (VAEs) (Mansimov et al., 2015) in 2015 and Generative Adversarial Networks (GANs) (Reed et al., 2016) in 2016 were capable of generating “visually plausible” images from text descriptions. Following these developments, diffusion models (Ho et al., 2020), such as DALLE-2 (OpenAI, 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022), emerged in 2022 and gained significant attention for their improvements in generating high-resolution, photorealistic images. Recently, even more powerful diffusion models, such as DALLE-3 (OpenAI, 2023) and Stable Diffusion 3 (AI, 2024a), have been released and are being used more widely in real-world applications (GOP, 2023; Davenport, 2023; Postma, [n. d.]). Despite these advancements in improving the quality of T2I generation, it remains uncertain whether these generative models perpetuate complex bias.
@@ 3.1. Detectors Selection and Implementation
We selected seven gender bias detectors, six of which have been widely used in previous studies to uncover gender bias in T2I models (Bansal et al., 2022; Cho et al., 2023; Karkkainen and Joo, 2021; Lee et al., 2024; Seshadri et al., 2023; Wang et al., 2024). We categorize these bias detectors into three classes: Vision-Language Models (e.g., CLIP (Radford et al., 2021), BLIP2 (Li et al., 2023)); Gender Classification Models, which are models specifically designed to classify the gender of a person from images (e.g., FairFace (Karkkainen and Joo, 2021), MiVOLO (Kuprashevich and Tolstykh, 2023)); and API Service (e.g., Face++ (Megvii, 2024)). A more detailed discussion is in Section 6 (Related Work). We include two detectors based on CLIP, designed to filter out images where gender is challenging to infer: CLIP-Prob (Seshadri et al., 2023) and CLIP-Uncertain (Bansal et al., 2022). Details of evaluated detectors are explained below:
@@
CLIP-Prob. This detector, proposed by Seshadri et al. (Seshadri et al., 2023), uses CLIP to predict gender and filters out low-confidence images. It employs a face detector [^5] https://ai.google.dev/edge/mediapipe/solutions/vision/face_detector/python to filter out images without any faces. Then, it also excludes images with a CLIP-predicted gender similarity (confidence level of the CLIP model) below 90%.
@@
3.2. Measuring Gender Bias in T2I Models
(1) Model Bias ${}\text{Model Bias Score}=\frac{1}{|P|}\sum_{p\in P}\frac{|n_{m}-n_{f}|}{(n_{m}+n_{f})}$
In the above equation, $N$ is the number of images generated for a given prompt, and $G_{i}$ represents the gender of the individual in image $i$. Otherwise, e.g., in the case of a low-quality image, $B_{i}$ is set to 0. $N_{\text{clear}}$ is the number of clear images (i.e., images that are not low-quality) generated for the prompt. For example, if a prompt generates 20 images and all 20 are male, the bias score is 1 (i.e., $(1\times 20)/20$). If eight images are male and 12 are female, the score is -0.2 (i.e., $(8-12)/20$). A score closer to +1 indicates that the model tends to generate more male images, and a score closer to -1 indicates that the model tends to generate more female images.
3.3. Dataset of Images Generated by T2I Models
(1) Prompt Generation. We describe a person using five dimensions: profession, personality, activity, object, and place. These dimensions combine categories discussed in existing literature (Cho et al., 2023; Bansal et al., 2022; Lin et al., 2023; Seshadri et al., 2023; Naik and Nushi, 2023; Wang et al., 2024; Friedrich et al., 2023). We compile words from previous above studies, removing duplicates and non-gender-neutral terms (e.g., actress). This resulted in 100 words: 40 for profession, 30 for personality, 10 for activity, 10 for object, and 10 for place. We choose not to scale up the word list as the human efforts to label these images are significant. We encourage interested researchers to extend the word list in future studies.
We use the template to generate prompts: “a photo of one real person, + [placeholder].” The placeholder can represent any five categories (e.g., profession). In the template, “one real” is included to condition T2I models to generate images with only one person and to ensure the image is realistic rather than cartoonish or stick-figure-like. The detailed template is shown in Table 1. For example, when we want T2I models to generate an image of a lawyer, the prompt will be: “a photo of one real person who is a lawyer.”
Table 1. Prompts used in the FairT2I Benchmark. The prefix is “a photo of one real person.”

### v2

@@ Abstract.
Text-to-Image (T2I) models have recently gained significant attention due to their ability to generate high-quality images and are consequently used in a wide range of applications. However, there are concerns about the gender bias of these models. Previous studies have shown that T2I models can perpetuate or even amplify gender stereotypes when provided with neutral text prompts (e.g., ‘a photo of a CEO’ is often associates with male images, while ‘a photo of nurse’ is often associates with female images). Researchers have proposed automated gender bias uncovering detectors for T2I models, but a crucial gap exists: no existing work comprehensively compares the various detectors and understands how the gender bias detected by them deviates from the actual situation.
This study addresses this gap by validating previous gender bias detectors using a manually labeled dataset and comparing how the bias identified by various detectors deviates from the actual bias in T2I models, as verified by manual confirmation. We create a dataset consisting of 6,000 images generated from three cutting-edge T2I models, Stable Diffusion XL, Stable Diffusion 3, and Dreamlike Photoreal 2.0. During the human-labeling process, we find that all three T2I models generate a portion (12.48% on average) of low-quality images (e.g., generate images with no face present), where human annotators cannot determine the gender of the person.
Our analysis reveals that all three T2I models show a preference for generating male images, with SDXL being the most biased. Additionally, images generated using prompts containing professional descriptions (e.g., lawyer or doctor) show the most bias. We evaluate seven gender bias detectors and find that none fully capture the actual level of bias in T2I models, with some detectors overestimating bias by up to 26.95%. We further investigate the causes of inaccurate estimations, highlighting the limitations of detectors in dealing with low-quality images. Based on our findings, we propose an enhanced detector called CLIP-Enhance, which most accurately measures the gender bias in T2I models, with a difference of only 0.47%-1.23%, and most effectively filters out 82.91% of low-quality images.  [^1] This paper potentially contains offensive information for some groups. We have made our dataset and code publicly available.  [^2] https://doi.org/10.6084/m9.figshare.27377649.v1
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755748isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Despite the advanced capabilities of T2I models, concerns about gender bias that are potentially demonstrated by these models remain significant. Previous studies have shown that T2I models tend to associate males with high-paying jobs, such as CEO, lawyer, and doctor, while associating females with low-paying jobs, such as housekeeper and cashier (Cho et al., 2023; Nicoletti and Bass, 2023). Bloomberg further highlighted how T2I models can exacerbate gender bias (Nicoletti and Bass, 2023). For example, women made up only about 3% of the images generated by the T2I model (Stable Diffusion v1.5) for the keyword “judge,” despite the fact that 34% of US judges are women. As AI-generated images increasingly permeate daily lives (Name, 2024), the stereotypes they reinforce and the neglect of minority genders may deepen existing unfairness. Therefore, it is crucial to uncover gender bias in T2I.
@@
  • Detector Enhancement. We enhance the detector, achieving top performance in accurately reflecting the gender bias in T2I models with a difference from the ground truth of only 0.47%-1.23% and effectively filtering out 82.91% of low-quality images.
@@ 2.1. Text-to-Image Models
A T2I model takes a textual description as input and generates a corresponding image that matches the description as output. With the advancement of deep learning techniques, T2I generation models began to evolve in the mid-2010s. Initially, Variational Auto Encoders (VAEs) (Mansimov et al., 2015) in 2015 and Generative Adversarial Networks (GANs) (Reed et al., 2016) in 2016 were capable of generating “visually plausible” images from text descriptions. Following these developments, diffusion models (Ho et al., 2020), such as DALLE-2 (OpenAI, 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022), emerged in 2022 and gained significant attention for their improvements in generating high-resolution, photorealistic images. Recently, more powerful diffusion models, such as DALLE-3 (OpenAI, 2023) and Stable Diffusion 3 (AI, 2024a), have been released and are being used widely in real-world applications (GOP, 2023; Davenport, 2023; Postma, [n. d.]). Despite these advancements in improving the quality of T2I generation, it remains uncertain whether these models perpetuate complex bias.
@@ 3.1. Detectors Selection and Implementation
We selected seven gender bias detectors, six of which have been widely used in previous studies to uncover gender bias in T2I models (Bansal et al., 2022; Cho et al., 2023; Karkkainen and Joo, 2021; Lee et al., 2024; Seshadri et al., 2023; Wang et al., 2024). We categorize these bias detectors into three classes: Vision-Language Models (e.g., CLIP (Radford et al., 2021), BLIP2 (Li et al., 2023)); Gender Classification Models, which are models specifically designed to classify the gender of a person from images (e.g., FairFace (Karkkainen and Joo, 2021), MiVOLO (Kuprashevich and Tolstykh, 2023)); and API Service (e.g., Face++ (Megvii, 2024)). A more detailed discussion is in Section 6 (Related Work). We include two detectors based on CLIP, designed to filter out images where gender is challenging to infer: CLIP-Prob (Seshadri et al., 2023) and CLIP-Uncertain (Bansal et al., 2022). Details of the evaluated detectors are in the supplementary material.
3.2. Measuring Gender Bias in T2I Models
(1) $\text{Model Bias Score}=\frac{1}{|P|}\sum_{p\in P}\frac{|n_{m}-n_{f}|}{(n_{m}+n_{f})}$
In the above equation, $N$ is the number of images generated for a given prompt, and $G_{i}$ represents the gender of the individual in image $i$. Otherwise, e.g., in the case of a low-quality image, $B_{i}$ is set to 0. $N_{\text{clear}}$ is the number of clear images (i.e., images that are not low-quality) generated for the prompt. For example, if a prompt generates 20 images and all 20 are male, the bias score is 1 (i.e., (1×20)/20(1$\times$ 20)/20( 1 × 20 ) / 20). If eight images are male and 12 are female, the score is -0.2 (i.e., (812)/20(8-12)/20( 8 - 12 ) / 20). A score closer to +1 indicates that the model tends to generate more male images, and a score closer to -1 indicates that the model tends to generate more female images.
3.3. Dataset of Images Generated by T2I Models
(1) Prompt Generation. We describe a person using five dimensions: profession, personality, activity, object, and place. These dimensions combine categories discussed in existing literature (Cho et al., 2023; Bansal et al., 2022; Lin et al., 2023; Seshadri et al., 2023; Naik and Nushi, 2023; Wang et al., 2024; Friedrich et al., 2023). We compile words from previous above studies, removing duplicates and non-gender-neutral terms (e.g., actress). This resulted in 100 words: 40 for profession, 30 for personality, 10 for activity, 10 for object, and 10 for place. We choose not to scale up the word list as the human efforts to label these images are significant. We encourage interested researchers to extend the word list in future studies. We use the template to generate prompts: “a photo of one real person, + [placeholder].” The placeholder can represent any five categories (e.g., profession). In the template, “one real” is included to condition T2I models to generate images with only one person and to ensure the image is realistic rather than cartoonish or stick-figure-like. For example, when we want T2I models to generate an image of a lawyer, the prompt will be: “a photo of one real person who is a lawyer.”
[…]
(3) Manually Labeling Gender Information. The labeling process follows three steps: (1) Preparation: Three authors initially discuss the labeling standards based on a preliminary set of 150 randomly selected images. (2) Labeling: Following the guidelines, the dataset is labeled by two annotators. Cohen’s Kappa coefficient (McHugh, 2012) is 0.86, indicating a high level of agreement (Landis and Koch, 1977; Lyu et al., 2024). (3) Discussion: After completing the labeling independently, two annotators met to discuss and resolve any discrepancies.
It is evident that all three models tend to generate more male images than female images, with an average of 63.57% of generated images being male and only 24.18% being female. On average, 12.48% of images are considered low quality. These results highlight the necessity of filtering out low-quality images when analyzing bias.
3.4. Measuring Difference between Actual and Detected Bias
(3) $\text{Prompt Bias Score Difference}=\frac{\sum_{i=1}^{N}|\text{PBS}^{i}_{\text{Detector}}-\text{PBS}^{i}_{\text{Actual}}|}{N}$
3.5. Measuring Detector Performance
Filtering Process. To evaluate the performance of detectors in the filtering process, we first formulate this process as a binary classification problem. Given an image, the detector classifies it as either a clear image or a low-quality image. Since the goal of the detector is to retrieve clear images, we use precision, recall, and F1-score (F1) to reflect the detector’s effectiveness in identifying clear images (Lyu et al., 2023). Additionally, we employ the filter rate (true negative rate) to assess the detector’s ability to filter out low-quality images, indicating how accurately the detector in this step identifies low-quality images. Specifically, to evaluate the detector’s performance in identifying clear images, we use precision, recall, and the F1. The F1 represents the harmonic mean of precision and recall. The formulas for these metrics are given below (P and R represent precision and recall, respectively):
[…]
(5) $\text{Filter Rate}=\frac{TN}{TN+FP}$
(6) $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
4. Results
4.1. RQ1: Bias in T2I Models
For RQ1, we present the gender bias in the T2I models. Table 6 (in supplementary material) displays the model bias score for SDXL, SD3, and Dreamlike. SDXL shows the highest bias (0.752), followed by SD3 (0.730), while Dreamlike is least biased (0.631). When analyzing prompt categories, profession is the most biased prompt category across all models, with SDXL (0.907) and SD3 (0.861) showing especially high scores. Using 100 prompts across three T2I models, we generate a total of 300 outputs. The distribution of prompt bias scores for all T2I models is illustrated in Figure 3 (in supplementary material). Specifically, 31.7% of outputs have a prompt bias score of 1, meaning the T2I model exclusively generates male images. Furthermore, 74.7% of outputs have a prompt bias score greater than 0, indicating a tendency towards generating male images. Nearly 90% of outputs have a prompt bias score either greater than 0.2 or less than -0.2, suggesting a strong tendency towards generating images of a single gender. These findings demonstrate that the current models exhibit a significant gender bias. Note that the detailed discussion of the RQ1 results is provided in the supplementary material.
Model SDXL SD3 Dreamlike
Ground Truth 0.752 0.730 0.631
CLIP 0.686 0.701 0.600
[…]
CLIP-Prob 0.820 0.794 0.801
[…]
CLIP-Uncertain 0.813 0.764 0.718
[…]
BLIP-2 0.796 0.755 0.662
[…]
Face++ 0.768 - 0.594
[…]
MiVOLO 0.759 0.716 0.613
[…]
FairFace 0.763 0.726 0.596
[…]
CLIP-Enhance 0.756 0.721 0.628
[…]
Model Bias Score for each T2I model. Table 1 shows the difference between different detectors and the ground truth. MiVOLO performs closest to the actual bias, with minimal percentage differences: 0.93% (SDXL), -1.92% (SD3), and -2.85% (Dreamlike). In contrast, CLIP-Prob introduces the largest deviations, overestimating the bias by 9.04% (SDXL), 8.77% (SD3), and 26.95% (Dreamlike). This discrepancy even changes the ranking of bias among models. For instance, CLIP-Prob incorrectly ranks Dreamlike as the second most biased model. CLIP-Uncertain also shows notable inaccuracies, with percentage differences: 8.11% (SDXL), 4.66% (SD3), and 13.80% (Dreamlike), all higher than the ground truth. FairFace performs as the second-best detector, with differences of 1.46% (SDXL), -0.55% (SD3), and -5.55% (Dreamlike). Face++ is unable to make predictions on images generated by SD3, the newest T2I model. BLIP-2 tends to overestimate bias across models, with percentage differences of 5.85% (SDXL), 3.42% (SD3), and 4.91% (Dreamlike). Overall, BLIP-2, CLIP-Prob, and CLIP-Uncertain overestimate gender bias, while CLIP, MiVOLO, and FairFace tend to underestimate gender bias.
Gender Model SDXL SD3 Dreamlike
CLIP 0.137 0.098 0.150
CLIP-Prob 0.503 0.526 0.443
CLIP-Uncertain 0.257 0.389 0.532
BLIP-2 0.245 0.188 0.298
Face++ 0.112 - 0.175
MiVOLO 0.158 0.073 0.209
FairFace 0.092 0.053 0.108
CLIP-Enhance 0.065 0.048 0.073
To understand why the seven detectors inaccurately detect gender bias in T2I models, we analyze two key components of the bias evaluation process described in Subsection 3.5. Specifically, we analyze the filtering process (precision, recall, F1-score, and filter rate) and the classification process (accuracy) using five metrics. The result is shown in Table 3. Detailed analysis of the filtering and classification processes is in the supplementary material.
Detector Filtering Process Accuracy (%)
Precision Recall F1 Filter Rate Male Female Overall
CLIP (Radford et al., 2021) 87.52 100.0 93.34 0.00 97.45 90.15 95.45
BLIP-2 (Li et al., 2023) 87.86 63.70 73.86 38.32 98.82 88.33 96.56
Face++ (Megvii, 2024) 95.03 97.94 96.46 71.24 98.12 77.06 91.58
CLIP-Enhance 97.56 97.54 97.55 82.91 98.95 92.38 97.13
5. Discussion
5.1. Enhancement Detector
(1) Face Detection: We use dlib’s face detection model (Dlib, 2017), as implemented in FairFace (Karkkainen and Joo, 2021), to filter out images without a clear face; (2) Multiple Person Filtering: We use YOLOv8 (Jocher et al., 2023) to exclude images containing multiple people. This is done by comparing the areas of the largest and second-largest bounding boxes (i.e., the rectangular outlines drawn around individuals within an image). If the second-largest box exceeds 50% of the largest, indicating that the presence of the second person may influence the perception of the main subject, the image is filtered out; (3) Cropping: We use YOLOv8 (Jocher et al., 2023) to draw bounding boxes around detected individuals and then crop the image based on these bounding boxes, ensuring that CLIP focuses on a single person.
Precision, Recall, F1, Filter Rate, and Accuracy. Table 3 presents these results. As discussed in Section 4.3, balanced detectors tend to perform better at bias detection. Our proposed detector, CLIP-Enhance, demonstrates a well-rounded performance across all metrics. CLIP-Enhance achieves the highest F1-score of 97.55, indicating its effectiveness in identifying clear images. When comparing to FairFace, which best filters low-quality images (although CLIP-Prob has a filter rate of 98.13% but a low recall of 19.06%), CLIP-Enhance achieves a filter rate of 82.91%, representing a 16% improvement over the baseline detector. Additionally, CLIP-Enhance achieves an overall accuracy of 97.13%, the second-highest.
5.2. Ethical Considerations
6. Related Work
T2I models have rapidly advanced in the past three years, drawing growing attention to bias in their generation process. When evaluating bias in T2I generation models, most bias evaluation metrics are classification-based, where characteristics are directly inferred. For studies using classification-based metrics, some rely on human-annotated gender in generated images for evaluation (Bansal et al., 2022; Fraser et al., 2023a, b; Wan and Chang, 2024). Bansal and Wan (Bansal et al., 2022; Wan and Chang, 2024) discuss the limitations of human annotation (e.g., limited budget for annotators from Amazon MTurk) and highlight the need for an effective automated annotation pipeline. The majority of studies using classification-based metrics rely on either classifier-based classification orvisual question answering (VQA)-based classification, both of which are automatic pipelines (Wan et al., 2024).
For classifier-based classification, several studies have used CLIP as a zero-shot classifier to determine the gender of generated images (Bansal et al., 2022; Seshadri et al., 2023; Orgad et al., 2023; Zhang et al., 2023; Lin et al., 2023; Lee et al., 2024; Kim et al., 2023). Bansal et al. (Bansal et al., 2022) is an early work that discusses gender bias in T2I models. They used the CLIP model for gender recognition. The FairFace classifier (Karkkainen and Joo, 2021), a pre-trained classifier on the FairFace dataset, has also been used to annotate gender (Friedrich et al., 2023, 2024). Naik et al. (Naik and Nushi, 2023) use Microsoft Cognitive Services as a classifier to determine gender. Wang et al. (Wang et al., 2024) further employ a metamorphic testing framework with Face++ (Megvii, 2024) to automatically reveal bias in image generation models. For VQA-based classification, Luccioni et al. (Luccioni et al., 2024) use BLIP (Li et al., 2022) to detect gender in generated images. Cho et al. (Cho et al., 2023), Esposito et al. (Esposito et al., 2023), and Wan et al. (Wan and Chang, 2024) use BLIP-2 (Li et al., 2023) for gender classification. Cho et al. (Cho et al., 2023) use an automatic framework to evaluate the bias in T2I generation models.
Fairness in Machine Learning (ML) extends beyond image-based systems to critical applications like Natural Language Processing (NLP), credit assessment, and healthcare. It has attracted attention from both the SE (Chen et al., 2023) and AI (Mehrabi et al., 2021) communities. Brun et al. (Brun and Meliou, 2018) called for SE research into fairness, while Zhang et al. (Zhang et al., 2020) framed it as a non-functional property. Chen et al. (Chen et al., 2024) surveyed fairness testing methods. Several studies have highlighted fairness challenges across domains. In NLP Systems, fairness issues in tasks like sentiment analysis and named entity recognition often arise from biased training data or model design, with various testing and mitigation methods proposed (Yang et al., 2021a; Mehrabi et al., 2020; Yang et al., 2021b; Wang et al., 2023; Yang et al., 2024b; Sze Khoo et al., 2023; D´ıaz et al., 2018; Liu et al., 2019; Asyrofi et al., 2021; Soremekun et al., 2022; Berk et al., 2021; Mehrabi et al., 2021; Chen et al., 2023; Guo et al., 2023). In Computer Vision Systems, studies have assessed fairness in deep image classification and benchmarked fairness-enhancing methods (Zhang et al., 2021a; Yang et al., 2024a). In Speech Recognition Systems, testing frameworks and input generation approaches have been developed (Rajan et al., 2022; Zhang et al., 2021b; Xiao et al., 2023; Lau et al., 2023).
7. Threats to Validity
8. Conclusion
This paper addresses the challenge of accurately detecting gender bias in Text-to-Image (T2I) models, highlighting discrepancies between biases identified by automated detectors and those verified through manual labeling. Our comprehensive study shows that existing detectors often fail to accurately reflect actual gender bias, with some overestimating bias by 27%. Key issues include low recall among high-accuracy detectors in gender classification and ineffective filtering of low-quality images by vision-language model-based detectors. To overcome these challenges, we propose an enhanced detector that accurately reflects gender bias in T2I models.
9. Acknowledgments
This research is supported by the Ministry of Education, Singapore under its Academic Research Fund Tier 3 (Award ID: MOET32020-0004). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendix A Detectors Selection and Implementation
We present the details of selection and implementation of the seven detectors used in this study:
@@ Appendix A Detectors Selection and Implementation
CLIP-Prob. This detector, proposed by Seshadri et al. (Seshadri et al., 2023), uses CLIP to predict gender and filters out low-confidence images. It employs a face detector (Google, 2024) to filter out images without any faces. Then, it also excludes images with a CLIP-predicted gender similarity (confidence level of the CLIP model) below 90%.
@@
Appendix B Dataset of Images Generated by T2I Models
B.1. Templates for Prompt Generation
The templates used to generate the prompts for the T2I models are shown in Table 4.
Table 4. Templates to generate the prompts. The prefix is “a photo of one real person.”